# A Comprehensive Benchmarking Analysis of Fault Recovery in Stream Processing Frameworks

### v1

@@ Abstract.
Nowadays, several software systems rely on stream processing architectures to deliver scalable performance and handle large volumes of data in near real time. Stream processing frameworks facilitate scalable computing by distributing the application’s execution across multiple machines. Despite performance being extensively studied, the measurement of fault tolerance—a key and most appealing feature offered by stream processing frameworks—has still not been measured properly with updated and comprehensive testbeds. Moreover, the impact that fault recovery can have on performance is mostly ignored. This paper provides a comprehensive analysis of fault recovery performance, stability, and recovery time in a cloud-native environment with modern open-source frameworks, namely Flink, Kafka Streams, and Spark Structured Streaming. Our benchmarking analysis is inspired by chaos engineering to inject failures. Generally, our results indicate that much has changed compared to previous studies on fault recovery in distributed stream processing. In particular, the results indicate that Flink can be the fastest and stablest under failures. Moreover, Kafka Streams shows performance instabilities after failures, which is due to its current repartitioning strategy that can be suboptimal in terms of load balancing. Spark Structured Streaming shows suitable fault recovery performance and stability, but with higher event latency. Our study intends to (i) help industry practitioners in choosing the most suitable stream processing framework for efficient and reliable executions of data-intensive applications; (ii) support researchers in applying and extending our research method as well as our benchmark; (iii) identify, prevent, and assist in solving potential issues in production deployments.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software performanceccs: Computer systems organization Cloud computingccs: Information systems Stream management
@@ 1. Introduction
Stream processing refers to a computing paradigm and architecture pattern for building distributed, event-based software systems that process massive amounts of data in near real time (Hirzel et al., 2014). Today’s stream processing paradigms enable software engineers to model their applications’ business logic as high-level representations in a directed acyclic graph (DAG), without explicitly defining the physical execution plan. This high level of abstraction is provided by industry-grade, open-source stream processing frameworks such as Spark (Armbrust et al., 2018), Flink (Carbone et al., 2015), or Kafka Streams (Wang et al., 2021). Such frameworks support software engineers in building highly scalable and efficient applications that process continuous data streams of massive volume. They provide high-level APIs and domain-specific languages to define the processing logic as directed acyclic processing graphs that filter, transform, aggregate, and merge data streams.
Over the last decade, several works have been published that evaluate the performance of distributed stream processing frameworks (Vogel et al., 2023; Henning and Hasselbring, 2024; Henning et al., 2024) or optimizations (Hirzel et al., 2014; Herodotou et al., 2020; Vogel et al., 2022). Beyond measuring performance metrics such as throughput and latency, fault tolerance stands as a critical requirement for continuously-operating production systems. Failures can occur unpredictably across various levels, from physical infrastructure to software layers. Stream processing systems, designed for continuous, low-latency processing, demand swift recovery mechanisms to tolerate and mitigate failures effectively.
@@
  • A comprehensive methodology for measuring fault recovery. The methodology covers stream processing’s most relevant metrics (Vogel et al., 2023), new metrics to improve fault recovery evaluations, a method for failure injection, and a method for automating fault recovery time measurements.
@@
  • An experimental analysis covering application metrics, system resources utilization, and fault recovery. [^1] On publication of this study, we will also provide as supplemental material an extended replication package of ShuffleBench for fault tolerance measurements. Moreover, the experimental data is also to be shared such that other researchers and practitioners may repeat and extend our work. Our evaluation covers the open-source stream processing frameworks Flink, Kafka Streams, and Spark Structured Streaming due to their industry acceptance and academic relevance (Vogel et al., 2023).
@@ 2.1. Distributed Stream Processing
Stream processing frameworks perform operations such as filterings, transformations, or aggregations in near real-time on continuous streams of data (Hirzel et al., 2014). State-of-the-art frameworks are designed for high throughput and low-latency processing, while also scaling with massive amounts of data (Fragkoulis et al., 2024; Henning and Hasselbring, 2024). To address these requirements, they run in a distributed fashion on commodity hardware. A key advantage of stream processing frameworks is that they provide dataflow models that abstract aspects such as cluster management, state management, and time semantics from their users (Sax et al., 2018). With such models, engineers describe the processing logic in DAGs of processing operators. The frameworks allow the initiation of multiple worker instances across various compute nodes with multiple threads, each instance handling a distinct portion of the data.
@@ 2.2. Fault Tolerance
In the absence of fault tolerance, stream processing applications would be forced to restart data processing from scratch whenever there is a loss due to a failure. Given that applications may have varying fault tolerance needs, the stream processing domain focuses on different processing semantics that precisely describe how a system is affected by failures (Fragkoulis et al., 2024). The relevant terms to characterize the fault tolerance semantics guarantees are at-most-once semantics (AMOS), at-least-once semantics (ALOS), and exactly-once semantics (EOS). Currently, the majority of stream processing applications demand reliable executions with strong semantics guarantees. Thus, AMOS is obsolete (Fragkoulis et al., 2024) and we focus on ALOS and EOS.
@@
In this paper, we focus on semantics guarantees regarding the system output because this is the most relevant for the majority of use cases and also covers the guarantees in the internal state updates. The stream processing frameworks provide different recovery semantic guarantees. This is important because the required guarantee varies according to the stream processing application and its use cases. In some scenarios, guaranteeing exactly-once can be of paramount importance, e.g., avoiding duplicate alerts in an observability platform. In other cases, at-least-once with rarely duplicated outputs is not equally detrimental, e.g. a monitoring system that receives twice the same CPU metric of a timestamp. We intend to measure the impacts of the stream processing frameworks when running with the fault recovery guarantees. Importantly, correct fault recovery is a crucial component of fault tolerance. While fault recovery focuses on restoring normal operation after a failure, fault tolerance aims to prevent downtime for high availability.
@@ 3.1. State-of-the-art
Recent publications have presented various experimental evaluations of fault recovery in stream processing (Vogel et al., 2023). Although many studies consider failures of worker instances (Lu et al., 2014; Qian et al., 2016; Lopez et al., 2016), the most updated and relevant is (van Dongen and van den Poel, 2021). It dives into the implementation, performance, and efficiency of fault recovery in four state-of-the-art stream processing frameworks: Spark Streaming, Flink, Spark Structured Streaming, and Kafka Streams. In addition, they evaluated the behavior of these frameworks under different types of faults and settings, including master failure with and without high-availability setups, driver failures for Spark frameworks, worker failure with or without exactly-once semantics, and application failures.
@@
Wang et al. (2021) explain Kafka Streams’ processing logic and demonstrate how Kafka Streams behaves in practice with large-scale deployments and performance insights exhibiting its flexible and low-overhead trade-offs. A very recent paper from Siachamis et al. (2024) focuses on complementary aspects of fault tolerance. The authors explored three approaches for checkpoint protocols (coordinated, uncoordinated, and communication-induced) aimed at fault tolerance in stream processing. The analysis delved into the theoretical strengths and weaknesses inherent in each approach. Their findings not only validated the prevailing preference for the coordinated approach but also illustrated instances, where the uncoordinated approach exhibited resilience and better performance.
Considering that many optimizations can be done for fault tolerance (Hirzel et al., 2014; Vogel et al., 2022), Su and Zhou (2021) address the issue of recovering from correlated failures in large-scale clusters. The authors propose an incremental and query-centric recovery paradigm. They formulate the problem of recovery scheduling under correlated failures and design algorithms to optimize the recovery latency.
@@ 3.2. Research Gaps
While fault recovery in stream processing has been addressed in various related studies, our understanding is that the methods and insights provided offer limited coverage. Despite being recognized as essential within a leading cloud observability platform, we identified significant research and analysis gaps in the state-of-the-art regarding fault recovery (Vogel et al., 2023):
  • Considering the continuous development of fault tolerance features and optimizations by the frameworks’ open-source communities, there is a need for updated measurements on frameworks’ executions facing worker instances’ failures.
  • The few studies that compare the frameworks regarding worker instances’ failures cover a single failure (van Dongen and van den Poel, 2021; Wang et al., 2022). However, a large physical machine crash impacts several smaller virtual machines or pods running within it. What can happen is that more than one worker instance fails. However, it remains unclear what is the impact of such a more critical failure in the running stream processing applications.
  • The existing literature only considers that worker instances can fail a single time (van Dongen and van den Poel, 2021; Wang et al., 2022). In real-world industry deployments, failures happen recurrently and at unpredictable times. One can assume that the frameworks’ recovery will always be the same, but we observed that such an assumption can have weak coverage with the reality of large deployments. This occurs mostly because the impact that a given failure has in the applications varies according to the exact moment that the failure happens (e.g., how far from the last saved state checkpoint), what working instances are affected by the failure, and the effectiveness of the subsequent recovery (Geldenhuys et al., 2022). Therefore, the level of fault recovery stability across recurrent failures remains unclear.
@@ 4.1. Benchmark
ShuffleBench utilizes and extends the benchmarking framework Theodolite (Henning and Hasselbring, 2022) to automate the benchmark execution in Kubernetes-based cloud environments. Figure 1 depicts our dataflow architecture for a corresponding stream processing application. In short, data records are read from a messaging system (Kafka in our case), assigned to keys, and shuffled such that all records having the same key are forwarded to the same worker instance. There, a key-specific custom aggregation is performed and, optionally, an output event is created that is then written back to Kafka. We refer to our previous work for further details about ShuffleBench (Henning et al., 2024).
@@
ShuffleBench has a flexible architecture and supports several parametric configurations. In this work, we extend ShuffleBench for fault tolerance measurements. Figure 2 provides an overview of ShuffleBench’s benchmark components and their extension to fault tolerance, where the scenarios and metrics are described in Section 5. We conduct our experimental evaluation in a Kubernetes cluster managed by the Elastic Kubernetes Service of Amazon cloud with the following pods running in cloud instances: 2 load generators (m6i.xlarge instances), 2 Kafka brokers (m6i.2xlarge), a manager node used on Flink and Spark (m6i.large), and 8 pods for worker instances with 1 CPU, 3 GB of RAM (running on c5.large instances).
In our experiments, we set up 20 000 real-time consumers that all have the same selectivity, which sum up to 50 %, meaning that each record is forwarded to $0.5$ consumers on average. Each consumer emits an output event for every record received. To consume and output results, 40 input and output Kafka partitions were used to maintain a suitable parallelism level that minimizes overheads. We run each experiment for 56 minutes in such a way that the benchmarks run for long enough to extract meaningful quantitative data, test with recurrent failures injected, and have sufficient time for every subsequent recovery.
@@ 4.2. Frameworks Deployment
We configured the frameworks on ShuffleBench deployment and parameters described in Section 4.1. To support fault tolerance, Flink (v. 1.17) and Spark Structured Streaming (v. 3.4) had checkpointing enabled for at-least-once semantics (ALOS), see more in LABEL:{sec:background-fault-tolerance}. Kafka Streams (v. 3.5) by default supports ALOS. Fault tolerance on Flink and Spark Structured Streaming involves periodic state backup via checkpointing, demanding a reliable filesystem to store state copies (van Dongen and van den Poel, 2021). We used in our setup an Amazon Elastic File System (EFS) filesystem volume mounted in the pods.
@@ 5.2. Failure Injection Scenarios
Experimental scenarios were inspired by chaos engineering (Basiri et al., 2016). Failures are injected using Chaos Mesh, [^3] https://github.com/chaos-mesh/chaos-mesh an open-source chaos engineering platform integrated with the Kubernetes deployment. We consider two scenarios:
@@
Random pods executing worker instances are deleted. When the pods are no longer accessible, both the records they were processing and their associated state are lost. Therefore, the frameworks have to trigger recovery from such a failure. Deployed with Kubernetes, new pods will be created and added to the pool of worker instances within an interval of seconds (usually taking from 2 to 10 seconds). We experimented with killing 1, 2, and 4 worker pods at once. Such values were determined to simulate simple and critical failure scenarios. Such failures are recurrently injected every 5 minutes to evaluate the fault recovery stability. This interval was defined for a reasonable recovery time between failures and also to avoid unnecessary long executions that would waste resources.
@@
Random worker instances pods face an operating system crash. This is expected to further test the resilience because the availability of resources is impacted. Contrasting with pods kill, in the pods crash failure no new pods are created to replace the failed ones. This failure also is expected to cause loss in the internal state of the worker instances affected, demanding failure recovery. Every failure impacts two random worker pods, which is a value set to represent a critical failure. The failures are recurrently injected every 10 minutes to evaluate the fault recovery stability and the failure remains for 5 minutes (value set according to feedback of industry partners). Then, executions can resume the normal state. Moreover, 5 minutes is a suitable value for creating and adding new virtual machines (Su and Zhou, 2021) to the pool.
@@ 5.3. Metrics
We collected several metrics of the benchmarking application. The metric to measure processing speed is throughput, measured in records processed per time unit and collected every second as a moving average of the last 5 seconds. A higher throughput is usually desired in many use cases. We measure input throughput refers to consuming from the Kafka input topic, and output throughput which is the application output written into another Kafka topic. We also measure event latency (the lower latency tends to be better) as an average of a 10-second window. We show latency percentiles to measure the distribution of events times: the p50 (50th percentile), p90, and p99. We also measure the consumer lag corresponding to the number of records buffered in the input topic.
The fault recovery on stream processing is measured with the metrics: Fault recovery performance, the impact of fault tolerance guarantees, stability, and fault recovery time (see Figure 2). We also present the following system monitoring indicators collected every 2 seconds from the cloud instances that run the worker pods: CPU utilization, memory utilization, and network traffic. [^4] To enhance visual clarity, the CPU utilization and network traffic of worker nodes are displayed using a 10-value moving window average.

### v2

@@ Abstract.
Nowadays, several software systems rely on stream processing architectures to deliver scalable performance and handle large volumes of data in near real-time. Stream processing frameworks facilitate scalable computing by distributing the application’s execution across multiple machines. Despite performance being extensively studied, the measurement of fault tolerance—a key feature offered by stream processing frameworks—has still not been measured properly with updated and comprehensive testbeds. Moreover, the impact that fault recovery can have on performance is mostly ignored. This paper provides a comprehensive analysis of fault recovery performance, stability, and recovery time in a cloud-native environment with modern open-source frameworks, namely Flink, Kafka Streams, and Spark Structured Streaming. Our benchmarking analysis is inspired by chaos engineering to inject failures. Generally, our results indicate that much has changed compared to previous studies on fault recovery in distributed stream processing. In particular, the results indicate that Flink is the most stable and has one of the best fault recovery. Moreover, Kafka Streams shows performance instabilities after failures, which is due to its current rebalancing strategy that can be suboptimal in terms of load balancing. Spark Structured Streaming shows suitable fault recovery performance and stability, but with higher event latency. Our study intends to (i) help industry practitioners in choosing the most suitable stream processing framework for efficient and reliable executions of data-intensive applications; (ii) support researchers in applying and extending our research method as well as our benchmark; (iii) identify, prevent, and assist in solving potential issues in production deployments.
@@ 1. Introduction
Stream processing refers to a computing paradigm and architecture pattern for building distributed, event-based software systems that process massive amounts of data in near real time (Hirzel et al., 2014). Today’s stream processing paradigms enable software engineers to model their applications’ business logic as high-level representations in directed acyclic graphs (DAGs), without explicitly defining the physical execution plan. This high level of abstraction is provided by industry-grade, open-source stream processing frameworks such as Spark (Armbrust et al., 2018), Flink (Carbone et al., 2015), or Kafka Streams (Wang et al., 2021). Such frameworks support software engineers in building highly scalable and efficient applications that process continuous data streams of massive volume. They provide high-level APIs and domain-specific languages to define the processing logic as directed acyclic processing graphs that filter, transform, aggregate, and merge data streams.
Over the last decade, several works have been published that evaluate the performance of stream processing frameworks (Vogel et al., 2023; Henning et al., 2024) or optimizations (Hirzel et al., 2014; Herodotou et al., 2020). Beyond measuring performance metrics such as throughput and latency, fault tolerance is a critical requirement for continuously operating production systems. Failures can occur unpredictably across various levels, from physical infrastructure to software layers. Stream processing systems, designed for continuous, low-latency processing, demand swift recovery mechanisms to tolerate and mitigate failures effectively.
@@
  • A comprehensive methodology for measuring fault recovery. The methodology covers stream processing’s most relevant metrics (throughput and latency) (Vogel et al., 2023) and new metrics to improve fault recovery measurements (fault recovery impact on performance, stability, and recovery time).
  • A method inspired by chaos engineering that enables the injection of different types of recurrent failures in stream processing experiments.
  • A method to automate fault recovery time measurements, available as supplementary material (Vogel et al., 2024b).
  • An experimental analysis covering application metrics, system resources utilization, and fault recovery. Our evaluation covers the open-source stream processing frameworks Flink, Kafka Streams, and Spark Structured Streaming due to their industry acceptance and academic relevance (Vogel et al., 2023).
The experimental data and an extended replication package of ShuffleBench (Henning et al., 2024) for fault tolerance measurements are available as supplementary material (Vogel et al., 2024b).
@@ 2.1. Distributed Stream Processing
Stream processing frameworks perform operations such as filterings, transformations, or aggregations in near real time on continuous streams of data (Hirzel et al., 2014). State-of-the-art frameworks are designed for high throughput and low-latency processing, while also scaling with massive amounts of data (Fragkoulis et al., 2024; Henning et al., 2024). To address these requirements, they run in a distributed fashion on commodity hardware, nowadays often in managed cloud environments.
A key advantage of stream processing frameworks is that they provide dataflow models that abstract aspects such as cluster management, state management, and time semantics from their users (Sax et al., 2018). The frameworks allow the initiation of multiple worker instances across various compute nodes with multiple threads, each instance handling a distinct portion of the data.
@@ 2.2. Fault Tolerance
In the absence of fault tolerance, stream processing applications would be forced to restart data processing from scratch whenever there is a loss due to a failure. Processing semantics guarantees describe how a system is affected by failures. Various guarantees are provided to accommodate the varying fault tolerance needs of different use cases (Fragkoulis et al., 2024). Relevant terms are at-most-once semantics (AMOS), at-least-once semantics (ALOS), and exactly-once semantics (EOS). Given that the majority of stream processing applications require reliable executions with strong semantics guarantees, we have shifted our focus from AMOS, which can be considered obsolete, to ALOS and EOS (Fragkoulis et al., 2024).
@@
In this paper, we focus on the semantic guarantees of the system output, as these are most relevant for the majority of use cases and also encompass the guarantees for internal state updates. The stream processing frameworks provide different recovery semantic guarantees. This is important because the required guarantee varies according to the stream processing application and its use cases.
In some scenarios, guaranteeing exactly-once delivery is of paramount importance, such as avoiding duplicate alerts in an observability platform. In other cases, at-least-once with rarely duplicated outputs is not equally detrimental, e.g., an information dashboard that receives twice the same CPU metric of a timestamp. We intend to measure the impacts of the stream processing frameworks when running with the fault recovery guarantees. Importantly, correct fault recovery is a crucial component of fault tolerance. While fault recovery focuses on restoring normal operation after a failure, fault tolerance aims to prevent downtime for high availability.
@@ 3.1. State-of-the-art
Recent publications provide experimental evaluations of fault recovery in stream processing (Vogel et al., 2023). Although many studies consider failures of worker instances (Lu et al., 2014; Qian et al., 2016; Lopez et al., 2016), the most updated and relevant is (van Dongen and van den Poel, 2021). It dives into the implementation, performance, and efficiency of fault recovery in four stream processing frameworks: Spark Streaming, Flink, Spark Structured Streaming, and Kafka Streams. In addition,  van Dongen and van den Poel (2021) evaluate the behavior of these frameworks under different types of faults and settings, including master failure with and without high-availability setups, driver failures for Spark frameworks, worker failure with or without exactly-once semantics, and application failures.
@@
Wang et al. (2021) demonstrate how Kafka Streams works on large-scale deployments and performance insights exhibiting its flexible and low-overhead trade-offs. A recent paper from Siachamis et al. (2024) focuses on complementary aspects of fault tolerance, exploring three approaches for checkpoint protocols (coordinated, uncoordinated, and communication-induced) aimed at fault tolerance in stream processing. The analysis delves into the approaches’ theoretical strengths and weaknesses. Their findings validate the preference for the coordinated approach and illustrated instances where the uncoordinated approach provides performance gains.
Considering that many optimizations can be done for fault tolerance (Hirzel et al., 2014), Su and Zhou (2021) address the issue of recovering from correlated failures in large-scale clusters. The authors propose an incremental and query-centric recovery paradigm. They formulate the problem of recovery scheduling under correlated failures and design algorithms to optimize the recovery latency.
@@ 3.2. Research Gaps
While fault recovery in stream processing has been addressed in various related studies, our understanding is that the methods and insights provided in the literature offer limited coverage. We believe that the following are relevant research gaps in the state-of-the-art regarding fault recovery on distributed stream processing:
  1. (1)
    Considering the continuous development of fault tolerance features and optimizations by the frameworks’ open-source communities (Fragkoulis et al., 2024), there is a need for updated measurements on frameworks’ executions facing worker instances’ failures.
  2. (2)
    Few studies compare how different frameworks recover worker instances and those that do focus on failures affecting a single instance (van Dongen and van den Poel, 2021; Wang et al., 2022). However, a large physical machine crash impacts several virtual machines or pods. What can happen is that more than one worker instance fails. However, the impact of such a more critical failure on running stream processing applications remains unclear.
  3. (3)
    The existing literature only covers scenarios where failures occur a single time (van Dongen and van den Poel, 2021; Wang et al., 2022). However, in real-world industry deployments, failures happen recurrently and unpredictably. One can assume that the frameworks’ recovery will always be the same, but we observed that such an assumption could have weak coverage with the reality of large deployments. This occurs mostly because the impact that a given failure has in the applications varies according to the exact moment that the failure happens (e.g., how far from the last saved state checkpoint), what working instances are affected by the failure, and the effectiveness of the subsequent recovery (Geldenhuys et al., 2022). Therefore, the level of fault recovery stability across recurrent failures demands more research.
@@ 4.1. Benchmark
ShuffleBench adopts and extends the benchmarking framework Theodolite (Henning and Hasselbring, 2022) to automate the benchmark execution in Kubernetes-based cloud environments. Figure 1 depicts our dataflow architecture for a corresponding stream processing application. Data records are read from a messaging system (e.g., Kafka), assigned to keys (zero to many), and shuffled such that all records having the same key are forwarded to the same worker instance. There, a key-specific custom aggregation is performed, and an output event is created that is then written back to Kafka. We refer to our previous work for a detailed description of ShuffleBench (Henning et al., 2024).
@@
ShuffleBench has a flexible architecture and supports several parametric configurations. In this work, we extend ShuffleBench for fault tolerance measurements. Figure 2 provides an overview of ShuffleBench’s components and their extension for fault tolerance (see the scenarios and metrics in Section 5). We conduct our experimental evaluation in a Kubernetes cluster managed by the Elastic Kubernetes Service of Amazon Web Services with the following pods running in dedicated cloud instances: 2 load generators (m6i.xlarge instances), 2 Kafka brokers (m6i.2xlarge), a manager node used on Flink and Spark (m6i.large), and 8 worker instances pods with 1 CPU, 3 GB of RAM (running on c5.large instances).
In our experiments, we set up 20 000 real-time consumers that all have the same selectivity, which sum up to 50 %, meaning that each record is forwarded to $0.50.5$ consumers on average. Each consumer emits an output event for every record received. To consume and output results, 40 input and output Kafka partitions were used to maintain a suitable parallelism level that minimizes overheads. We run each experiment for 56 minutes in such a way that the benchmarks run for long enough to extract meaningful quantitative data, test with recurrent failures injected, and have sufficient time for every subsequent recovery.
@@ 4.2. Frameworks Deployment
We configured the frameworks on ShuffleBench deployment and parameters described in Section 4.1. To support fault tolerance, Flink (v. 1.17) and Spark Structured Streaming (v. 3.4) had checkpointing enabled for at-least-once semantics (ALOS), see more in LABEL:{sec:background-fault-tolerance}. Kafka Streams (v. 3.5) by default supports ALOS. Fault tolerance on Flink and Spark Structured Streaming involves periodic state backup via checkpointing, demanding a reliable filesystem to store state copies (van Dongen and van den Poel, 2021). We used in our setup an Amazon Elastic File System (EFS) volume mounted in the pods.
@@ 5.2. Failure Injection Scenarios
Our experimental scenarios were inspired by chaos engineering (Basiri et al., 2016) using Chaos Mesh [^2] Chaos Mesh is an open-source platform: https://github.com/chaos-mesh/chaos-mesh. to inject failures in two scenarios:
@@
Random pods executing worker instances are deleted. When the pods are no longer accessible, both the records they were processing and their internal state are lost. Therefore, the frameworks trigger recovery from such a failure. Deployed with Kubernetes, new pods will be created and added to the pool of worker instances within an interval of seconds (usually taking from 2 to 10 seconds). We experimented with killing 1, 2, and 4 worker pods at once. Such values were determined to simulate simple and critical failure scenarios. The failures are recurrently injected every 5 minutes to evaluate the fault recovery stability. This interval was defined for a reasonable recovery time between failures and also to avoid unnecessary long executions that would waste resources.
@@
Random worker instances pods face an operating system crash. This is expected to further test the resilience because the availability of resources is impacted. Contrasting with pods kill, in the pods crash failure no new pods are created to replace the failed ones. This failure also is expected to cause loss in the internal state of the worker instances affected, demanding failure recovery. Every failure impacts two random worker pods, which is a value set to represent a critical failure. The failures are recurrently injected every 10 minutes to evaluate the fault recovery stability and the failure remains for 5 minutes (value set according to feedback of industry partners). Then, executions can resume the normal state. Moreover, 5 minutes is a suitable value for creating and adding new virtual machines to the pool (Su and Zhou, 2021).
@@ 5.3. Metrics
We collected several metrics of the benchmarking application. The metric to measure processing speed is throughput, measured in records processed per time unit and collected every second as a moving average of the last 5 seconds. A higher throughput is usually desired in many use cases. We measured input throughput (the rate of records consumed from the Kafka input topic, see Fig. 1) and output throughput (the rate of records written to the Kafka output topic). We also measure event latency (lower is better) as an average of a 10-second window. We show latency percentiles to measure the distribution of events times: the median (p50), 90th percentile (p90), and 99th percentile (p99). We also measure the consumer lag corresponding to the number of records buffered in the input topic.
The fault recovery on stream processing is measured with the metrics: Fault recovery performance, the impact of fault tolerance guarantees, stability, and fault recovery time (see Figure 2). We also present the following system monitoring indicators collected every 2 seconds from the cloud instances that run the worker pods: CPU usage, memory utilization, and network traffic. [^3] To enhance visual clarity, the CPU usage and network traffic of worker nodes are displayed using a 10-value moving window average.
The method described here is applied in Sections 6, 7 and 8 to conduct measurements and extract insights. In this paper, we show the most insightful results, complementary results can be found in the supplementary material (Vogel et al., 2024b).